brian is 2 years older than bob who is 7 years older than mark. if their combined age 61 years, find the age of each person
Answer by mananth(16949)

Mark =x
Bob = x+7
Brian = x+7+2
Total =61
x+x+7+x+9=61
3x+16=61
3x=61-16
3x=45
/3
x=15
Mark is 15 so you can get others ages